a dielectric material has relative permittivity 18 and loss tangent 10^-3 at 100 MHz find the conductivity of the medium

1
Expert's answer
2021-11-18T15:01:59-0500

"\sigma=w \epsilon_r\epsilon_0tan\phi"

"f=10^8hz\\\epsilon_r=18"

Put value


"\sigma=2\pi\times10^8\times18\times8.85\times10^{-12}\times10^{-3}=1.0\times10^{-4} mho /m"